OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). ?

OpenStudy (anonymous):

Write the equation of given line in the form of y = mx + b, where m is the slope. Perpendicular lines have slopes that when multiplied equal -1. So, you can find the slope of the perpendicular line that way. Then, find the equation of that line.

OpenStudy (radar):

Good advice, arrange the 2x-3y = 3 to look like y=mx+b known as the slope intercept form. If this is a problem, I am standing by to assist if needed.

OpenStudy (anonymous):

i did that and got -7.3 um im really confused -_-

OpenStudy (radar):

3y=2x-3 This is the begining y=(2/3)x -1 This is the given line in slope intercept for Note that the slope is 2/3 Now at GT pointed out your perpendicular line will a slope of a negative 1 (-1) divided by this slope or the negtative reciprocal so the slope of the perpendicular line is a -3/2 m=-3/2 did you follow with understanding?

OpenStudy (anonymous):

kinda

OpenStudy (radar):

O.K. we have the slope of the perpendicular line m=-3/2 We were also given a point (-8,2). We can now put the perpendicular line in the correct slope intercept form by simply plugging our now known values: 2=(-3/2)(-8)+b Note 2 is the Y, -3/2 is the slope m, -8 is the x

OpenStudy (radar):

Solving for b, 2=12+b b=10 We have it all y=(-3/2)x+10